OpenStudy (anonymous):

I need some help with Trigonometry? I'm scared of failing an upcoming test and would like to review one problem.

OpenStudy (anonymous):

I'm just starting out with this practice booklet and I'm worried I'm doing something wrong. This is the question: When Θ = 5 pi over 3, what are the reference angle and the sign values for sine, cosine, and tangent?

OpenStudy (anonymous):

you can find \(\frac{5\pi}{3}\) on the last page of the attached trig cheat sheet, that is most useful

OpenStudy (anonymous):

My answer choices are: Θ'= pi over 3, cosine is positive, sine and tangent are negative Θ'= negative pi over 3, sine and cosine are positive, tangent is negative Θ' = negative 5 pi over 3, cosine is positive, sine and tangent are negative Θ'= 5 pi over 3, sine and cosine are positive, tangent is negative I think it's A, however I'm also leaning towards D.

OpenStudy (anonymous):

i have to say i never really understood what a "reference angle" is, or what the purpose of it was, but i think it is the smallest angle that the terminal side makes with the x axis

OpenStudy (anonymous):

yeah i think it is A as well

OpenStudy (anonymous):

for one thing " cosine is positive, sine and tangent are negative" is true

OpenStudy (anonymous):

I don't either, and I wish my workbook would have gone further into detail with it instead of just saying "Quadrant A is all positive always." Quite annoying when you'd like to study.

OpenStudy (anonymous):

Yeah I was thinking the same thing. I'm going to go ahead and take this test and I'll let you know if it's correct, if you'd like. :)

OpenStudy (anonymous):

yeah, i just googled it it says "reference angle is always positive" and always in quadrant 1, i.e. always \(0\leq \theta \leq \frac{\pi}{2}\) i frankly think it is a waste of time

OpenStudy (anonymous):

you don't need no stinkin reference angle to find sine, cosine, tangent etc just know where the heck on the unit circle you are

OpenStudy (anonymous):

You aren't the only one! It wouldn't be so bad if it wasn't so complex, you know? If they would give you one thing (e.g. sine, etc.) to work with at a time, it would probably be much easier.

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

A was correct! You, my friend, are getting a medal from me. :) I can't tell you how much I appreciate you helping me!

mathslover (mathslover):

Reference angle for a standard angle refers to an angle made by the \(x\) axis and the terminal side of the given angle. Reference angle is always an acute angle. (Credits to SparkNotes for the image attached)